OpenStudy (anonymous):

Can someone show me the work for this problem? 25h+40=-15h-80

OpenStudy (wwwhhhaaattt?):

Hold on a sec

OpenStudy (wwwhhhaaattt?):

@StudyGurl14

JoelTheBoss (joel_the_boss):

Simplifying 25h + 40 = 15h + -80 Reorder the terms: 40 + 25h = 15h + -80 Reorder the terms: 40 + 25h = -80 + 15h Solving 40 + 25h = -80 + 15h Solving for variable 'h'. Move all terms containing h to the left, all other terms to the right. Add '-15h' to each side of the equation. 40 + 25h + -15h = -80 + 15h + -15h Combine like terms: 25h + -15h = 10h 40 + 10h = -80 + 15h + -15h Combine like terms: 15h + -15h = 0 40 + 10h = -80 + 0 40 + 10h = -80 Add '-40' to each side of the equation. 40 + -40 + 10h = -80 + -40 Combine like terms: 40 + -40 = 0 0 + 10h = -80 + -40 10h = -80 + -40 Combine like terms: -80 + -40 = -120 10h = -120 Divide each side by '10'. h = -12 Simplifying h = -12

OpenStudy (anonymous):

25h + 40 = -15 h - 80 25h + 15h = -80 - 40 40 h = -120 h = -3

OpenStudy (studygurl14):

h = -3

OpenStudy (mathstudent55):

@Joel_the_boss You copied the problem incorrectly. You left out a negative sign on the right side.